# Modeling Language as a Sequence of Thoughts

### v1

@@ 5 Conclusion
Appendix A Memory Gate Dynamics Over Training
Each TG cross-attention block (i.e., the layers that read from the sentence-memory) includes a learnable scalar memory gate for that scales the memory cross-attention update before it is added through the residual path. Thus, corresponds to an unscaled memory contribution, while amplifies the influence of the memory pathway relative to the rest of the layer’s residual computation. Because the self-attention residual branch is not scaled by this parameter, the gate provides a direct readout of the model’s relative weighting of sentence-memory content at each depth.
@@ Appendix A Memory Gate Dynamics Over Training
Appendix B Numerical Results
B.1 Scaling Efficiency Results
@@
B.2 GPT-2 with Sentence Boundary Bias Results
| Training Tokens | Standard GPT-2 | GPT-2 + Sent. Boundary | Standard TG |
@@ B.2 GPT-2 with Sentence Boundary Bias Results
B.3 TG with Fixed Token-Span Recurrence Results
| Training Tokens | Standard TG | TG + 25-Span | TG + 50-Span | TG + 75-Span | Standard GPT-2 |
@@ B.3 TG with Fixed Token-Span Recurrence Results
B.4 GPT-2 + Gist Masking Results
| Training Tokens | Standard GPT-2 | GPT-2 + Gist Masking | Standard TG |
@@ B.4 GPT-2 + Gist Masking Results
Appendix C Other TG Design Principles
C.1 Batch construction: Uniform Token-Budget Bucketing
C.2 Scheduled Dropouts
In TG apply stochastic regularization in several places.

### v2

@@ 5 Conclusion
In conclusion, this work shows considerable promise for the use of latent thought gestalts learned end-to-end via next-token prediction training objective. The critical next step is to scale this approach to fully realize its potential.
Appendix A Recursive Backpropagation Through Working Memory
TG writes sentence vectors to working memory without detaching their computation graphs. While the memory size limits which sentence vectors are available to cross-attention in the forward pass, it does not limit the backward flow of gradients. The loss at step backpropagates recursively through the stored sentence vectors into older sentences that influenced the computation of the current memory entries. This recursive chain terminates at the first sentence of the stream (Sentence 1), which is computed with an empty memory. Thus, the maximum backward depth is bounded by the maximum stream length (§2.1).
Appendix B Other TG Design Principles
B.1 Batch construction: Uniform Token-Budget Bucketing
B.2 Stochastic Regularization (Scheduled Dropouts)
In TG, we apply stochastic regularization in several places.
Appendix C Memory Gate Dynamics Over Training
Each TG cross-attention layer (i.e., the layers that read from the sentence-memory) includes a learnable scalar memory gate for that scales the memory cross-attention update before it is added through the residual path. Thus, corresponds to an unscaled memory contribution, while amplifies the influence of the memory pathway relative to the rest of the layer’s residual computation. Because the self-attention residual branch is not scaled by this parameter, the gate provides a direct readout of the model’s relative weighting of sentence-memory content at each depth.
@@ Appendix C Memory Gate Dynamics Over Training
Appendix D Component Gradient Share
Figure 8 shows a systematic shift in where the optimization signal concentrates as training data increases. From M to M to M tokens, the average gradient share assigned to sentence-memory cross-attention rises (6%7%9%), while the share assigned to within-sentence self-attention decreases (34%33%30%); the feedforward/MLP components remain approximately constant (60–61%). We interpret the increasing cross-attention share as evidence that, at larger pretraining scales, the learned sentence gestalts become more informative, and TG correspondingly allocates more learning capacity to the working-memory pathway for next-token prediction. This trend is consistent with the learned memory-gate dynamics (Appendix C) and with the strong degradation observed when gradient flow through memory is removed (§3.5).
Appendix E Numerical Results
E.1 Scaling Efficiency Results
@@
E.2 GPT-2 with Sentence Boundary Bias Results
| Training Tokens | GPT-2 | GPT-2 + Sent. Boundary | Standard TG |
@@ E.2 GPT-2 with Sentence Boundary Bias Results
E.3 TG with Fixed Token-Span Recurrence Results
| Training Tokens | Standard TG | TG + 25-Span | TG + 50-Span | TG + 75-Span | GPT-2 |
@@ E.3 TG with Fixed Token-Span Recurrence Results
E.4 GPT-2 + Gist Masking Results
| Training Tokens | GPT-2 | GPT-2 + Gist Masking | Standard TG |